OpenStudy (anonymous):

We are given two numbers and we know that twice the smaller one is four plus the larger one. We also know that their sum is 23. What are the numbers? In working with my tutor, I came to the conclusion that x = 9 and y = 14. but I can't decipher how we got there? help?

OpenStudy (anonymous):

\[2x=4+y\] \[x+y=23\] solve simultaneously

OpenStudy (anonymous):

9 times 2 = 18 subtract 4 = 14 SOOOO therefore..9 IS the smaller number..and 14 is the larger number ^w^ and then you add the 9 and 14 together to get 23 :p

OpenStudy (anonymous):

\[2x=4+23-x\]

OpenStudy (anonymous):

3x=27 x=9